Here is my problem as written in my homework:
A metallurgist needs to produce 10kg of an alloy which is 65% copper and 35% tin. He has an alloy which is 50% copper and 50% tin. How much of this 50-50 alloy must he melt together with pure copper to make the desired alloy? How much pure copper must he use?
Here is how I attempted to solve the problem:
(10-x)(.50)(.50)+(x)(1)=10(.65)(.35)
So the answer that I got was that he would need 2.03 kg of the pure copper and 7.97 kilograms of the 50-50 mixture.
When I substituted the numbers into my equation, it did not come out right.
Thank you for your time and patience.

Answer by lwsshak3(11628)

Here is my problem as written in my homework:
A metallurgist needs to produce 10kg of an alloy which is 65% copper and 35% tin. He has an alloy which is 50% copper and 50% tin. How much of this 50-50 alloy must he melt together with pure copper to make the desired alloy? How much pure copper must he use?
**
let x=amt of pure copper to be melted with 50-50 alloy
10-x=amt of 50-50 alloy to be melted with pure copper
..
100%*x+50%*(10-x)=65%*10
x+(10-x)/2=6.5
2x+10-x=2*6.5=13
x+10=13
x=3
10-x=7
amt of pure copper to be melted with 50-50 alloy=3kg
amt of 50-50 alloy to be melted with pure copper=7kg